What is the rule for x 1 2 3 4 5 6
y-1-4-7-10-13-16

Found 2 solutions by stanbon, wgomero:
Answer by stanbon(75887)

Since there is a constant difference of -3 in the y-values
there is a linear relation.
You have two points (1,-1), (2,-4)
Solving for the slope and the y-intercept you should get
y=-3x+2
Cheers,
Stan H.

Answer by wgomero(28)

Resolution:
as the x number increase the y decrease by 3.
1 = -1
2 = -1 -3 = -4
3 = -4 -3 = -7
4 = -7 -3 = -10
5 = -10 -3 = -13
6 = -13 - 3 = -16
7 = -16 -3 = -19